What is the sum of the digits of the four-digit numbers formed using the digits
2,4,6,8(repetitions not allowed).

Found 2 solutions by khwang, trip.3c:
Answer by khwang(438)

Sum of 2,4,6,8 is 20.
There are 4!=24 such four-digit numbers.
So, the sum of them is 480.
Somehow, this question is not quite clear and doesn't make much sense.
Kennny

Answer by trip.3c(1)

Total no of arrangements possible = 4! = 24
So, no of occurences for all the 4 digits unit place = 24/4 = 6times. So sum of all such unit numbers = 6 [ 8 + 6 + 4 + 2 ] = 6*20 = 120
Similarly, for tenth, hundredth and thousandth place
The sum of 24 times 120 in 4 places(thousandth, hundredth, tenth and unit) = 133320